OpenStudy (anonymous):

Use the table to find each probability. P(counselor junior) Table: ----------------------- |Grade Level| Male| Female| ----------------------- |Junior | 18 | 21 | ----------------------- |Senior | 25 | 16 | -----------------------

OpenStudy (anonymous):

Sorry title of the table is Characteristics of Comp Counselors. if that's needed

OpenStudy (kirbykirby):

If you're just looking for the probability that a counsellor is a junior, then they can be male or female. So.. All juniors = 18+21 = 39 Now, you have the probability space being all juniors + seniors, since a counsellor can be a junior or a senior, which is = 18 + 21 + 25 + 16 = 80 Thus, P(counsellor junior) = 39/80